OpenStudy (anonymous):

Solve t = the quantity p plus 3 r, divided by six for r.

OpenStudy (anonymous):

\[t=\frac{p+3r}{6}\]?

OpenStudy (anonymous):

\[6t=p+3r\] \[6t-p=3r\] \[\frac{6t-p}{3}=r\]